# A countable-support symmetric iteration separating $\mathsf{PP}$ from $\mathsf{AC}$

### v5

@@ 5.3 The modified limit filter
The companion paper [5] defines the limit-stage filter as the normal closure of head pullbacks. In this manuscript we work with the strengthened filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50) obtained by additionally adjoining the subgroups $\Delta_{\lambda}^{\uparrow}(E,D)$ for globally admissible pairs $(E,D)$.
@@
Theorem 5.59 ($\mathcal{M}\models\mathsf{ZF}$).
Let $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic and $G_{\lambda}:=G_{\infty}\cap\mathbb{P}_{\lambda}$. Let $\operatorname{HS}_{\mathrm{Ord}}$ be the class of hereditarily symmetric (cf. Definition 5.56) $\mathbb{P}_{\mathrm{Ord}}$-names (class-names) with respect to $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$, and define
@@
Work in the background metatheory used to formalize class-length iterations (cf. [5, §(Class-length iterations)] and Remark 5.60). In this setting, definable class recursion suffices to define the valuation map for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names and to carry out the usual recursion defining hereditarily symmetric class-names.
For each set ordinal $\lambda$, the restriction to stage $\lambda$ yields the set-length symmetric system $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ and hence $M_{\lambda}\models\mathsf{ZF}$ by Theorem 5.58. Moreover, every set-sized $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{x}$ is supported in some stage $\lambda$, so $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\lambda}}\in M_{\lambda}$. Consequently,
$\mathcal{M}=\{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda},$
so $\mathcal{M}$ is a transitive class.
The verification of the $\mathsf{ZF}$ axioms in $\mathcal{M}$ follows by the standard class-length symmetric iteration argument (ibid.): each axiom instance with set parameters from $\mathcal{M}$ is witnessed at some stage $\lambda$ and hence holds in $M_{\lambda}$, and therefore holds in $\mathcal{M}$. ∎
@@ 5.4 Bookkeeping: names versus interpretations
Remark 5.60 (Background metatheory for class-length constructions).
To speak formally about an $\mathrm{Ord}$-length bookkeeping function and an $\mathrm{Ord}$-length symmetric iteration as class objects, we work in a background theory of classes supporting definable class recursion for the iteration, e.g. $\mathsf{GBC}+\mathsf{ETR}$, and we fix a class well-order of $V$ (equivalently, assume Global Choice in the background). Its set universe is denoted $V$ and satisfies $\mathsf{ZFC}$. All symmetric models $\mathcal{N}$, intermediate stages $M_{\alpha}$, and the final model $\mathcal{M}$ are constructed inside this set universe; only the bookkeeping map $\mathcal{B}$ and the iteration itself are treated as class objects in the ambient background.
@@
Let $f\in\mathcal{M}$ be a surjection instance of one of our templates (PP or $\mathsf{AC}_{\mathsf{WO}}$). Choose $\dot{f}\in\operatorname{HS}$ with $f=\dot{f}^{G}$, and let $\beta:=\sup(I(\dot{f}))+1$ as in Lemma 5.67, so $f\in M_{\beta}$. Then there exists some $\alpha\geq\beta$ such that the bookkeeping schedule hits $(\beta,c_{\dot{f}})$ at stage $\alpha$, and hence:
  • either $f$ already splits in $M_{\alpha}$, in which case $f$ splits in $\mathcal{M}$;
  • or $f$ does not split in $M_{\alpha}$, and the successor-stage forcing at $\alpha$ uses the corresponding orbit package, so $f$ splits in $M_{\alpha+1}$ and hence in $\mathcal{M}$.
@@
By Definition 5.62 (unboundedness of the enumeration), there is some $\alpha\geq\beta$ with $\mathcal{B}(\alpha)=(\beta,c_{\dot{f}})$. At that stage, we run the bookkeeping test on $\dot{f}$ in $M_{\alpha}$. If it succeeds, the orbit package forcing at stage $\alpha$ adds a splitting object, which persists to all later stages. If it fails, then $f$ already has a splitting object in $M_{\alpha}$, and this also persists. In either case, $f$ splits in $\mathcal{M}$. ∎
@@
5.5 The symmetric iteration (black box)
We now invoke the iteration framework via Remark 2.4. Working in the background metatheory, we form a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$. Concretely, we build an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §3.6], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$. The next subsection states the properties of the final symmetric model that we prove.
@@ 5.5 The symmetric iteration (black box)
At successor stages, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the modified stage filters from Definition 5.50. Concretely, $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pullbacks $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$, together with admissible subgroups $\Delta_{\alpha+1}^{\uparrow}(E,D)$. In particular, for each active package coordinate $d$ at stage $\alpha+1$ whose countable support is witnessed by $E_{d}$, the generator $\Delta_{\alpha+1}^{\uparrow}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.47 and 5.48.
At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group as provided by the symmetric-iteration framework (Remark 2.4(a),(b)). The limit-stage filter is $\tilde{\mathscr{F}}^{*}_{\lambda}$ as defined in Definition 5.50, and its normality and $\omega_{1}$-completeness are recorded in Lemma 5.55. For set-localization in the $\mathrm{Ord}$-length model we use only the stage-support lemma (Lemma 5.67) below. We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@ 5.6 What the iteration yields
Lemma 5.67 (Stage-localization of set-names at class length).
@@
where $\operatorname{supp}(p)\subseteq\mathrm{Ord}$ is the (countable) support of the condition $p$. Then $I(\dot{x})$ is a set of ordinals, hence $\alpha:=\sup(I(\dot{x}))+1<\mathrm{Ord}$. Moreover:
@@
    $\dot{x}$ is in fact a $\mathbb{P}_{\alpha}$-name (all conditions appearing in $\dot{x}$ lie in $\mathbb{P}_{\alpha}$),
@@
    for any $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ and $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, we have $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
@@
Define $I(\dot{x})$ by recursion on the rank of the name $\dot{x}$. Since $\dot{x}$ is set-sized, Replacement and Union yield that $I(\dot{x})$ is a set of ordinals, hence bounded in $\mathrm{Ord}$. Let $\alpha=\sup(I(\dot{x}))+1$. Every condition $p$ appearing in $\dot{x}$ has $\operatorname{supp}(p)\subseteq I(\dot{x})\subseteq\alpha$, so $p\in\mathbb{P}_{\alpha}$, hence $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name. The valuation recursion for $\dot{x}^{G_{\infty}}$ consults only conditions appearing in $\dot{x}$, all of which lie in $\mathbb{P}_{\alpha}$, hence depends only on $G_{\alpha}$. Therefore $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$. ∎
@@
Choose a hereditarily symmetric set-name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$. By Lemma 5.67, letting $\alpha:=\sup(I(\dot{f}))+1$, we may view $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name and we have
$f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}\qquad\text{and}\qquad\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}.$
Thus $f=f_{\alpha}\in M_{\alpha}\subseteq\mathcal{M}$.
If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done. Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping. By the final paragraph of Definition 5.62, there is some stage $\beta\geq\alpha$ such that either $f_{\alpha}$ already splits in $M_{\beta}$ (in which case we are done), or else the successor iterand at stage $\beta$ is the orbit $\mathsf{PP}$-package $\mathbb{Q}_{[f_{\alpha}]}$.
Assume we are in the latter case, and fix such a stage $\beta$. Let $\dot{s}_{f_{\alpha}}$ be the canonical $\mathbb{Q}_{[f_{\alpha}]}$-name for the generic section in the $f_{\alpha}$-coordinate. In the full package extension, this coordinate generic yields a section of $f_{\alpha}$ (cf. Proposition 4.15).
By Lemma 5.47, $\dot{s}_{f_{\alpha}}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f_{\alpha}]},\mathscr{G}_{[f_{\alpha}]},\mathscr{F}_{[f_{\alpha}]})$ over $M_{\beta}$. By Lemma 5.49, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence
$s_{f_{\alpha}}:=\dot{s}_{f_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}.$
Since $s_{f_{\alpha}}$ is a section of $f_{\alpha}$, it remains a section of $f$ in the final model $\mathcal{M}$. Therefore every $f:Y\twoheadrightarrow X\subseteq T$ in $\mathcal{M}$ splits, i.e. $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$. ∎
@@
At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support cofinal in $\lambda$, so the usual stage-bounding arguments do not apply. In this manuscript we handle such limits by using the modified limit filter (Definition 5.50), whose normality and $\omega_{1}$-completeness are recorded in Lemma 5.55, together with the core admissibility lemma (Lemma 5.53). This matches the treatment of countable-cofinality limits in the general iteration framework ([5, §3.5–§3.6]).

### v6

@@ 5.3 The modified limit filter
The companion paper [5] defines the limit-stage filter from head pullbacks alone: at uncountable cofinality as the normal filter generated by those pullbacks, and at countable cofinality as the smallest normal $\omega_{1}$-complete filter containing them. In this manuscript we work with the strengthened filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.53) obtained by additionally adjoining the subgroups $\Delta_{\lambda}^{\uparrow}(E,D)$ for globally admissible pairs $(E,D)$.
@@
Lemma 5.62 (Upward persistence of hereditarily symmetric names).
Let $\beta\leq\lambda$ be stages, where $\lambda$ may also be $\mathrm{Ord}$. If $\tau\in\operatorname{HS}_{\beta}$, then, viewed as a $\mathbb{P}_{\lambda}$-name via canonical inclusion, $\tau\in\operatorname{HS}_{\lambda}$.
Proof.
We argue by induction on the rank of $\tau$.
Since $\tau\in\operatorname{HS}_{\beta}$, every subname $\sigma$ of $\tau$ belongs to $\operatorname{HS}_{\beta}$. By the induction hypothesis, each such $\sigma$ also belongs to $\operatorname{HS}_{\lambda}$.
It remains to prove that $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$. Because $\tau\in\operatorname{HS}_{\beta}$, we have
$\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\beta}.$
By Definition 5.53, the pullback
$(\pi^{\beta}_{\lambda})^{-1}\!\big[\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\big]$
belongs to $\tilde{\mathscr{F}}^{*}_{\lambda}$.
We claim that
$(\pi^{\beta}_{\lambda})^{-1}\!\big[\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\big]\subseteq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau).$
Indeed, if $k\in\mathscr{G}_{\lambda}$ satisfies $\pi^{\beta}_{\lambda}(k)\in\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)$, then by Lemma 5.47,
$k\cdot\tau=\pi^{\beta}_{\lambda}(k)\cdot\tau=\tau.$
So $k\in\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)$, proving the claim.
By upward closure of $\tilde{\mathscr{F}}^{*}_{\lambda}$, it follows that $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$. Therefore $\tau\in\operatorname{HS}_{\lambda}$. ∎
Remark 5.63 (Why the standard symmetric-extension proof applies at length $\mathrm{Ord}$).
Although $\mathbb{P}_{\mathrm{Ord}}$ is a class forcing, every $\mathbb{P}_{\mathrm{Ord}}$-name used in the final model construction is a set. The usual proof of the symmetric extension theorem (for example, [8, Lemma 15.51]) is a recursion on the ranks of set-sized names and uses only:
  1. (i)
    the valuation map for set-sized names,
  2. (ii)
    the action of the automorphism group on names, and
  3. (iii)
    that the filter of subgroups is proper and normal.
In the present setting these ingredients are available for $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$: valuation is defined in the background metatheory from Remark 5.65, the group action is defined stagewise by the iteration construction, and properness/normality are provided by Lemma 5.58. Thus the same proof applies to the final class of hereditarily symmetric names.
Theorem 5.64 (The final class hereditarily symmetric model).
Let $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic. Let $\operatorname{HS}_{\mathrm{Ord}}$ be the class of hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names (cf. Definition 5.59) with respect to $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$, and define
@@
Work in the background metatheory from Remark 5.65. By Remark 5.63, the standard proof of the symmetric extension theorem applies to the class system $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$. Hence the class of interpretations of hereditarily symmetric names,
$\mathcal{M}=\{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\},$
satisfies $\mathsf{ZF}$.
For transitivity, let $x\in\mathcal{M}$ and $y\in x$. Choose $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ with $x=\dot{x}^{G_{\infty}}$. Then there is some pair $\langle\dot{y},p\rangle\in\dot{x}$ with $p\in G_{\infty}$ and $y=\dot{y}^{G_{\infty}}$. Since $\dot{x}$ is hereditarily symmetric, $\dot{y}\in\operatorname{HS}_{\mathrm{Ord}}$, so $y\in\mathcal{M}$. ∎
@@ 5.4 Bookkeeping: names versus interpretations
Remark 5.65 (Metatheory for bookkeeping).
We continue with the metatheoretic conventions of Remark 2.1. For the $\mathrm{Ord}$-length bookkeeping in Section 5.4 (in particular, to define the class function $\mathcal{B}:\mathrm{Ord}\to V$ in Definition 5.67), we work in a background two-sorted structure $(V,\mathcal{C})\models\mathsf{GBC}+\mathsf{ETR}$ and, only for this bookkeeping/canonical-enumeration purpose, fix a class well-order of $V$ (equivalently, assume Global Choice in the background). All forcing semantics for $\mathbb{P}_{\mathrm{Ord}}$ are understood externally exactly as in Remark 2.1 (in particular, we do not assume a global forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$).
@@
Let $f\in\mathcal{M}$ be a surjection instance of one of our two templates ($\mathsf{PP}$ or $\mathsf{AC}_{\mathsf{WO}}$). Choose a set-sized $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ with $f=\dot{f}^{G_{\infty}}$. By Lemma 5.73, there is $\beta<\mathrm{Ord}$ such that $\dot{f}$, viewed as a $\mathbb{P}_{\beta}$-name, belongs to $\operatorname{HS}_{\beta}$ and
$f=\dot{f}^{G_{\beta}}\in M_{\beta}.$
Then there exists some $\alpha\geq\beta$ such that the bookkeeping schedule hits $(\beta,c_{\dot{f}})$ at stage $\alpha$, and at that stage:
  • either $f$ already has the required splitting object in $M_{\alpha}$ (a section in the $\mathsf{PP}$-case, a right inverse in the $\mathsf{AC}_{\mathsf{WO}}$-case), or
  • the successor-stage forcing at $\alpha$ uses the corresponding orbit package.
Consequently, $f$ has the required splitting object in $\mathcal{M}$.
@@
By the unboundedness clause in Definition 5.67, there exists $\alpha\geq\beta$ with $\mathcal{B}(\alpha)=(\beta,c_{\dot{f}})$.
View $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name via canonical inclusion. By Lemma 5.62, this included name belongs to $\operatorname{HS}_{\alpha}$. Since $\dot{f}$ is already a $\mathbb{P}_{\beta}$-name, its interpretation at stage $\alpha$ is still $f$:
$\dot{f}^{G_{\alpha}}=\dot{f}^{G_{\beta}}=f.$
Thus, when the bookkeeping test is run on $(\beta,c_{\dot{f}})$ at stage $\alpha$, clauses (i), (ii), and (iii) of Definition 5.67 all hold exactly when $f$ is a genuine stage-$\alpha$ instance of the required template. If $f$ already has the required splitting object in $M_{\alpha}$, we are done.
Otherwise the bookkeeping test succeeds, so the stage-$\alpha$ iterand is the corresponding orbit package $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$. In the package extension this adds the required splitting object for $f$, and by the hereditary-symmetry lifting lemmas used at successor stages, that object belongs to $M_{\alpha+1}\subseteq\mathcal{M}$. Hence $f$ has the required splitting object in $\mathcal{M}$. ∎
@@
5.5 The $\mathrm{Ord}$-length symmetric iteration
We now carry out the specific $\mathrm{Ord}$-length recursion used in this paper. Working in the background metatheory from Remark 2.1, we define by transfinite recursion on $\alpha\in\mathrm{Ord}$ an iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, starting from $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and producing intermediate symmetric models $M_{\alpha}$. At successor stages we use the orbit-package iterands described below; at limit stages we use the countable-support forcing limit and the corresponding group/projection infrastructure from the bounded-stage framework, together with the modified limit filter from Definition 5.53. No general class-length theorem from [5] is invoked at this point; only the bounded-stage infrastructure summarized in Remark 2.4 is used. The next subsection states the properties of the final symmetric model that we prove.
@@ 5.5 The $\mathrm{Ord}$-length symmetric iteration
At successor stages, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the modified stage filters from Definition 5.53. Concretely, $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pullbacks $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$, together with admissible subgroups $\Delta_{\alpha+1}^{\uparrow}(E,D)$. In particular, for each active package coordinate $d$ at stage $\alpha+1$ whose countable support is witnessed by $E_{d}$, the generator $\Delta_{\alpha+1}^{\uparrow}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.50 and 5.51.
At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group as provided by the symmetric-iteration framework (Remark 2.4(a),(b)). The limit-stage filter is $\tilde{\mathscr{F}}^{*}_{\lambda}$ as defined in Definition 5.53, and its normality and $\omega_{1}$-completeness are recorded in Lemma 5.58. For set-localization in the $\mathrm{Ord}$-length model we use the forcing-localization lemma (Lemma 5.72) together with the descent lemma for set-sized final hereditarily symmetric names (Lemma 5.73) below. We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@ 5.6 What the iteration yields
Lemma 5.72 (Forcing-localization of set-sized class-length names).
@@
where $\operatorname{supp}(p)\subseteq\mathrm{Ord}$ is the (countable) support of the condition $p$. Then $I(\dot{x})$ is a set of ordinals, hence
$\alpha:=\sup(I(\dot{x}))+1<\mathrm{Ord}.$
Moreover:
@@
    every subname of $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name, and in particular $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name;
@@
    for any $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ and $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, one has $\dot{y}^{G_{\infty}}=\dot{y}^{G_{\alpha}}$ for every subname $\dot{y}$ of $\dot{x}$, and in particular $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
@@
Define $I(\dot{x})$ by recursion on the rank of the name $\dot{x}$. Since $\dot{x}$ is set-sized, Replacement and Union imply that $I(\dot{x})$ is a set of ordinals, hence bounded in $\mathrm{Ord}$. Let $\alpha=\sup(I(\dot{x}))+1$.
If $\dot{y}$ is a subname of $\dot{x}$, then by construction $I(\dot{y})\subseteq I(\dot{x})\subseteq\alpha$. Hence every condition appearing in $\dot{y}$ has support contained in $\alpha$, so $\dot{y}$ is a $\mathbb{P}_{\alpha}$-name. This proves (1).
For (2), fix a subname $\dot{y}$ of $\dot{x}$. Since all conditions appearing in $\dot{y}$ lie in $\mathbb{P}_{\alpha}$, the valuation recursion for $\dot{y}^{G_{\infty}}$ consults only $G_{\alpha}$, so $\dot{y}^{G_{\infty}}=\dot{y}^{G_{\alpha}}$. The special case $\dot{y}=\dot{x}$ gives the final statement. ∎
Lemma 5.73 (Stage descent of set-sized hereditarily symmetric names).
Let $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ be set-sized. Then there exists an ordinal $\alpha<\mathrm{Ord}$ such that:
  1. 1.
    every subname of $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name;
  2. 2.
    every subname of $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$;
  3. 3.
    for every $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$, with $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, one has
    $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}.$
In particular, $\dot{x}^{G_{\infty}}\in M_{\alpha}$.
Proof.
For each subname $\dot{y}\in\operatorname{tc}(\dot{x})\cup\{\dot{x}\}$, since $\dot{y}\in\operatorname{HS}_{\mathrm{Ord}}$, we have
$\operatorname{Stab}_{\mathscr{G}_{\mathrm{Ord}}}(\dot{y})\in\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}}.$
By Lemma 5.60, choose a countable globally admissible pair $(E_{\dot{y}},D_{\dot{y}})$ at stage $\mathrm{Ord}$ such that
$\Delta_{\mathrm{Ord}}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})\subseteq\operatorname{Stab}_{\mathscr{G}_{\mathrm{Ord}}}(\dot{y}).$
For each such $\dot{y}$, let
$S(\dot{y}):=\{\xi<\mathrm{Ord}:\exists z\ \langle\xi,z\rangle\in D_{\dot{y}}\}.$
Since $D_{\dot{y}}$ is countable and consists of package coordinates, $S(\dot{y})$ is a countable set of ordinals.
Define
$J(\dot{x}):=I(\dot{x})\cup\bigcup_{\dot{y}\in\operatorname{tc}(\dot{x})\cup\{\dot{x}\}}\bigl(E_{\dot{y}}\cup S(\dot{y})\bigr).$
Because $\dot{x}$ is set-sized, the collection $\operatorname{tc}(\dot{x})\cup\{\dot{x}\}$ is a set, and each $E_{\dot{y}}\cup S(\dot{y})$ is countable. Hence $J(\dot{x})$ is a set of ordinals. Let
$\alpha:=\sup(J(\dot{x}))+1<\mathrm{Ord}.$
By Lemma 5.72, every subname of $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name, and the valuation of every subname depends only on $G_{\alpha}$. This proves (1) and (3), once we show that every subname belongs to $\operatorname{HS}_{\alpha}$.
We prove by induction on rank that every subname $\dot{y}$ of $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$. So fix such a $\dot{y}$, and assume inductively that every proper subname of $\dot{y}$ belongs to $\operatorname{HS}_{\alpha}$.
Because $E_{\dot{y}}\subseteq J(\dot{x})\subseteq\alpha$, and every coordinate in $D_{\dot{y}}$ has stage index in $S(\dot{y})\subseteq J(\dot{x})\subseteq\alpha$, we have $D_{\dot{y}}\subseteq\mathrm{Coords}_{\alpha}$. Moreover, the defining admissibility condition $\operatorname{Fix}(E_{\dot{y}})\leq\operatorname{Stab}_{\mathscr{G}}(d)$ for $d\in D_{\dot{y}}$ is independent of the ambient stage once the coordinates lie below that stage. Hence $(E_{\dot{y}},D_{\dot{y}})$ is globally admissible at stage $\alpha$.
Let
$k\in\Delta_{\alpha}^{\uparrow}(E_{\dot{y}},D_{\dot{y}}).$
By Corollary 5.46, there exists
$\hat{k}\in\Delta_{\mathrm{Ord}}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})$
with $\pi^{\alpha}_{\mathrm{Ord}}(\hat{k})=k$. Since
$\Delta_{\mathrm{Ord}}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})\subseteq\operatorname{Stab}_{\mathscr{G}_{\mathrm{Ord}}}(\dot{y}),$
we have $\hat{k}\cdot\dot{y}=\dot{y}$. Now $\dot{y}$ is a $\mathbb{P}_{\alpha}$-name, so by Lemma 5.47,
$k\cdot\dot{y}=\pi^{\alpha}_{\mathrm{Ord}}(\hat{k})\cdot\dot{y}=\hat{k}\cdot\dot{y}=\dot{y}.$
Hence
$\Delta_{\alpha}^{\uparrow}(E_{\dot{y}},D_{\dot{y}})\subseteq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{y}).$
By Lemma 5.60, this implies $\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{y})\in\tilde{\mathscr{F}}^{*}_{\alpha}$.
Together with the induction hypothesis for the subnames of $\dot{y}$, this shows $\dot{y}\in\operatorname{HS}_{\alpha}$. The induction is complete, so (2) holds.
Finally, (3) is exactly the valuation clause from Lemma 5.72, and therefore
$\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}\in M_{\alpha}.$
Proposition 5.74 (Stage capture).
For every $x\in\mathcal{M}$, there exists $\alpha<\mathrm{Ord}$ such that $x\in M_{\alpha}$. Equivalently,
$\mathcal{M}=\bigcup_{\alpha\in\mathrm{Ord}}M_{\alpha}.$
Proof.
Write $x=\dot{x}^{G_{\infty}}$ for some set-sized $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$. By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$. Hence $x\in M_{\alpha}$. ∎
@@
Choose a set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ such that $f=\dot{f}^{\,G_{\infty}}$. By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{f}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and
$f=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}.$
If $f$ already has a section in $M_{\alpha}$, we are done.
By Lemma 5.68, there is a stage $\beta\geq\alpha$ such that the successor iterand at stage $\beta$ is the orbit $\mathsf{PP}$-package $\mathbb{Q}_{[f]}$. Fix such a stage $\beta$.
Let $\dot{s}_{f}$ be the canonical $\mathbb{Q}_{[f]}$-name for the generic section in the $f$-coordinate. In the full package extension, this coordinate generic yields a section of $f$ (cf. Proposition 4.15).
By Lemma 5.50, $\dot{s}_{f}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f]},\mathscr{G}_{[f]},\mathscr{F}_{[f]})$ over $M_{\beta}$. By Lemma 5.52, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence
$s_{f}:=\dot{s}_{f}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}.$
Since $s_{f}$ is a section of $f$, it remains a section of $f$ in the final model $\mathcal{M}$. Therefore every $f:Y\twoheadrightarrow X\subseteq T$ in $\mathcal{M}$ splits, i.e. $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$. ∎
@@
At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support cofinal in $\lambda$, so the usual stage-bounding arguments do not apply. In this manuscript we handle such limits by using the modified limit filter (Definition 5.53), whose normality and $\omega_{1}$-completeness are recorded in Lemma 5.58, together with the core admissibility lemma (Lemma 5.56). This matches the bounded-stage treatment of countable-cofinality limits in [5, Limit filter and Remark (Failure of stage-bounding at countable cofinality)].